OpenStudy (anonymous):

@Abhisar Please help me solve this question. A corner on a racing track has a radius of curvature of 125m. What is the minimum coefficient of static friction between the tyres of a car and the track so that a car does not skid when taking the bend at 108km/h?

OpenStudy (anonymous):

NO i have to write that again :(

OpenStudy (anonymous):

???

OpenStudy (anonymous):

 \[F _{c} = m \frac{ v ^{2} }{ r }\] \[friction .force = \mu mg\] u is the coefficient of friction and to prevent the car from skidding 2 forces must be at least equal. \[m \frac{ v ^{2} }{ r } = \mu mg\] \[\mu = \frac{ v ^{2} }{ r g }\]